# Real Noncommutative Convexity II: Extremality and nc convex functions

### v2

@@ 3. Complexification of extremals
and $\sum\,c(\alpha_{k})^{\mathsf{T}}c(\alpha_{k})=1_{2n}.$ Since $c(x)$ is pure in $K$ we have that a positive scalar multiple of $c(\alpha_{k})$ is real isometry matrix $\eta_{k}\in M_{2n_{k},2n}(\mathbb{R})$, with $\eta_{k}^{\mathsf{T}}c(x_{k})\eta_{k}=c(x)$. It is then easy to see (since $c$ is one-to-one and ‘modular’) that a positive scalar multiple of $\alpha_{k}$ is a complex isometry $\beta_{k}\in M_{n_{k},n}(\mathbb{C})$, with $\beta_{k}^{*}x_{k}\beta_{k}=x$. So $x$ is pure in $K_{c}$. If $c(x)$ is extreme in $K$ then in addition each $c(x_{k})$ commutes with $c(\alpha_{k})\,c(\alpha_{k})^{\mathsf{T}}$ and $c(x_{k})\,\eta_{k}\,\eta_{k}^{\mathsf{T}}=\eta_{k}c(x)\eta_{k}^{\mathsf{T}}$. So $x_{k}$ commutes with $\alpha_{k}\alpha_{k}^{*}$. The rest is easy. ∎
@@
The converse of the lemma is false even in the simplest example of $K$ singleton: $1$ is pure and extreme in $\mathrm{ncS}(\mathbb{C})$ but $c(1)=1\oplus 1$ is not pure, and is clearly not extreme. The class of extreme points considered in this lemma is interesting in some examples, but not in others.
@@
Remarks. (On real parts of complex matrix states.) 1) By the argument in [29, Proposition 5.3.3], every pure state $\varphi$ on a real operator system $V$ is the restriction to $V$ of the real part of a pure state on $V_{c}$ (not usually $\varphi_{c}$). However not every pure (resp. extreme) element $x\in K$ is the real part of a pure (resp. extreme) element of $K_{c}$, sadly. See [33], or 2) below, or Example 3.5. In the latter example, since $\ell^{\infty}_{2}$ has no irreducible representation into $M_{n}(\mathbb{C})$ for $n\geq 2$ we have no extreme points in level 2 (or higher) of the complexification. Similarly for pure points: in that example there are no pure elements of $K_{2}={\rm UCP}(\mathbb{C},M_{2}(\mathbb{C}))$ by 6.1.7 in [14].
@@
The classical argument that pure/extreme states are the real part of a pure extension to the complexification is a standard one in convexity (finding an extreme point of the set of extensions and showing that this is extreme in the superspace). That the nc extension of this fails illustrates the failure in the nc situation of this very common trick in convexity. We do not know a characterization of when a nc pure (resp. extreme) in K is the real part of some pure (resp. extreme) in $K_{c}.$
@@
2) (A quaternionic counterexample.) Similarly the identity map on $M_{2}(\mathbb{C})$ is extreme, so maximal. However its real part $\psi$, which is Re, restricted to the copy of the quaternions $\mathbb{H}$, is not real maximal, so not extreme. This can be seen from Proposition 2.12 (2), or by the following direct argument: $\psi$ is the compression to the first two rows and columns of the usual (irreducible) representation of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. So it is not maximal by Proposition 2.12 (2). It is interesting that $\psi$ is pure, by Proposition 2.12 (1). (Cf. [33, Example 3.27].)
@@
The latter also answers the question: For a real ucp map $\varphi$, there exists a maximal dilation of its complexification. Is the real part of this maximal dilation a maximal dilation of $\varphi$? In the example above, the complexification of $\psi$ has maximal dilation $x\mapsto x\oplus x$ for $x\in M_{2}(\mathbb{C})$. The real part of this is a real dilation of $\psi$. It is the map $\psi\oplus\psi$. However this is not maximal, since a summand of a maximal is easily seen to be maximal, and $\psi$ is not maximal. This also follows by Proposition 2.12 (2).
@@
For $K$ a complex compact nc convex set, if $x\in K$ is complex pure (resp. extreme, maximal) in $K$ then $x$ is real pure (resp. extreme, maximal) in $K$. Conversely, if $x$ is real extreme and complex-irreducible then $x$ is complex extreme and irreducible in $K$. In particular, in general
@@
Suppose that $x$ is real extreme in $K$. Then $x$ is real, and hence complex, maximal in $K$. A variant of the argument in Corollary 3.2 shows that if $x$ is irreducible for $M_{n}(\mathbb{C})$ then it is complex extreme. Finally we give an example where $\partial_{\mathbb{C}}K\neq\partial_{\mathbb{R}}K$ in the next remark. ∎
Remark. By a slight variant of the argument for the last assertion of Corollary 3.2, $x\in K_{1}$ is real extreme in $K_{1}$ if and only if $x$ is complex extreme in $K$. However generally the converses in Lemma 3.8 fail for pure (resp. extreme). E.g. if $x$ is real extreme in $K$ then $x$ need not be complex extreme in $K$. This fails even in $K_{2}$. To see this consider an extreme $x\in K$ such that $x$ is not complex nc extreme in $K_{c}$, such as in Example 3.5. However by the Remark after Example 3.5, $x$ is real nc extreme in $K_{c}$. Similarly $x$ is real pure in $K_{c}$, but not complex pure there (if it were it would be complex extreme since it is also complex maximal).
@@ 4. Noncommutative convex functions
In the complex case this is simply a remark in [14] without proof. We add a proof, partly to serve as a model for the later more complicated multivalued nc function case. For $f\in M_{n}(B(K))$, we have that $f$ is graded, and so $\text{Epi}(f)$ is graded. Also $f$ preserves direct sums, which results in $\text{Epi}(f)$ being closed under direct sums. Indeed, let $(x_{i},\beta_{i})\in\text{Epi}_{m_{i}}(f)$, and let $\alpha_{i}\in M_{k,m_{i}}(\mathbb{R})$ be a family of isometries such that $\sum_{i}\alpha_{i}\alpha_{i}^{\mathsf{T}}=I_{k}$. We need to check that
@@
Remark. Remark $7.2.2$ in [14] holds in the real case with the same proof. Similarly, the Hansen-Pedersen type application in [14, Proposition 7.2.3] holds in the real case, showing that $f$ is real nc convex if and only if it is classically convex at every ‘level’.
@@
Let $f\in M_{n}(B(K))$, then $f$ is real nc convex (resp. lowersemicontinuous) if and only if $f_{c}=\Psi(f+i0)$ is complex convex (resp. lowersemicontinuous) as an element of $M_{n}(B(K_{c}))$.
@@
Suppose that $f+ig\in M_{n}(B(K))_{c}=M_{n}(B(K_{c}))$ for a real compact nc convex set $K$. Then $f+ig$ is complex nc convex if and only if $c(f,g)$ is real nc convex in $M_{2n}(C(K))$. If these hold then $f$ is real nc convex.
@@
Now $(f+ig)(x+iy)=u_{m}^{*}(f(c(x,y))+ig(c(x,y)))u_{m}$. Since $u_{m}(I_{n}\otimes\alpha)=(I_{2n}\otimes\alpha)u_{m}$, it suffices to show that
@@
Now [14, Proposition 7.2.8] is provable by complexification:
@@ 5. Convex multivalued functions and convex envelopes
We now extend our real versions of convexity and lower semicontinuity to bounded multivalued selfadjoint nc functions $F$. We define the latter in the real case as in [14, Definition 7.3.1] except of course matrices have real entries (including unitary matrices, naturally). Thus e.g. $F$ is bounded if there is a constant $\lambda>0$ such that for every $\beta\in F(x)$ there is an $\alpha\in F(x)$ with $\alpha\leq\beta$ and $\|\alpha\|\leq\lambda$. The graph ${\rm Graph}(F)$ is defined by
@@
Convexity and lower semicontinuity of $F$ are defined just as in [14, Definition 7.3.2], except all matrices have real entries. That is, $F$ is convex (resp. lowersemicontinuous) if ${\rm Graph}(F)$ is nc convex (resp. closed). We say that $F\leq G$ if $G(x)\subseteq F(x)$ for all $x\in K$. Every real (bounded) nc function $f:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ gives a (bounded) multivalued selfadjoint nc function $F(x)=[f(x),\infty)$, with Epi $(f)={\rm Graph}(F)$. So $f$ is convex if and only if $F$ is convex.
@@
For a real nc convex set $K$, every real bounded nc multivalued selfadjoint function $F:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ has a unique complex bounded nc multivalued extension $F_{c}:K_{c}\to M_{n}(\mathcal{M}(\mathbb{C}))$, and this is selfadjoint. Similarly every real bounded nc selfadjoint multivalued function $F:K\to M_{n}(\mathcal{M}(\mathbb{C}))$ has a unique complex bounded selfadjoint nc multivalued extension $F_{c}:K_{c}\to M_{n}(\mathcal{M}(\mathbb{C}))$. Finally, $F\mapsto F_{c}$ is an order embedding: $F\leq G$ if and only if $F_{c}\leq G_{c}$.
@@
$F_{c}(\beta(x+iy)\beta^{*})$ $=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,F(c(\beta)c(x+iy)c(\beta)^{\mathsf{T}})\,\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,c(\beta)F(c(x+iy))c(\beta)^{\mathsf{T}}\,\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=\beta\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,F(c(x+iy))\,\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}\beta^{*}$
@@
which shows the unitary invariance of the extension. A similar proof shows the direct sum property. To see that $F_{c}(x+iy)=F_{c}(x+iy)+M_{n}(\mathcal{M}(\mathbb{C}))^{+}$ choose $\alpha\in M_{n}(\mathcal{M}(\mathbb{C}))^{+}$. Then $\alpha=u^{*}c(\alpha)u.$ For $z\in F(c(x+iy))$ we know that $z+c(\alpha)\in F(c(x+iy))$. Hence
@@
Thus $F_{c}(x+iy)=F_{c}(x+iy)+M_{n}(\mathcal{M}(\mathbb{C}))^{+}$.
@@
$(x+iy)\oplus(x-iy)=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}c(x+iy)\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}$
@@
$G(x+iy)\oplus G(x-iy)$ $=G\Big{(}\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}c(x+iy)\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}\Big{)}$
$=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}F(c(x+iy))\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}$.
Here we first used that $G(x\oplus y)=G(x)\oplus G(y)$. The last part of the above equation is a $2n\times 2n$ matrix with top-right and bottom-left corners being 0. Comparing the top-left corners of the matrices in the above equation gives
$G(x+iy)=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,F(c(x+iy))\,\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
as desired.
@@
so taking the real part of both sides, and using that Re is unitarily equivariant (see Lemma 2.5 in [8]), we get that $\pi(G)$ is unitarily equivariant. For direct sums, take $x,y\in K$ and then $G(x)\oplus G(y)=G(x\oplus y)$. Re respects direct sums so taking Re of both sides gives that $\pi(G)$ respects direct sums.
Fix $x\in K_{m}$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$. If $z+\alpha\in\pi(G)(x)+M_{n}(M_{m}(\mathbb{R}))^{+}$ for $z=\text{Re }(y)$ where $y\in G(x)$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$, then $y+\alpha\in G(x)$ and so $\text{Re }(y)+\alpha\in\pi(G)(x)$. Thus $\pi(G)$ is upward directed. To show boundedness, let $x\in K$ and $\beta\in\pi(G)(x)$ so that $\beta=\text{Re }y$ for some $y\in G(x+i0)$. Because $G$ is bounded there exists a $\lambda>0$ and $\alpha\in G(x+i0)$ with $\alpha\leq y$ and $\alpha||\leq\lambda$. Taking the Re of the inequality and using that Re is a contraction gives $\text{Re }\alpha\leq\text{Re }y=\beta$ and $\text{Re }\alpha||\leq\lambda$. Thus, $\lambda$ is a bound for the function $\pi(G)$.
@@
The converse follows from the lemma. Alternatively, suppose $F_{c}$ is convex and let $x\in K$ and $\alpha$ be an isometry. Then (2) holds for $F$ because $F(\alpha^{\mathsf{T}}x\alpha)=F_{c}(\alpha^{\mathsf{T}}x\alpha)$ and since (2) holds for $F_{c}$.
@@
The rest of [14, Section 7.3] is the same in the real case. Turning to [14, Section 7.4] we define the real convex envelope $\bar{F}$ of a real bounded selfadjoint multivalued nc function $F:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ to be the multivalued nc function $\bar{F}:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ with Graph $(\bar{F})=\overline{{\rm ncconv}}({\rm Graph}(F)).$ It is easy to see that [14, Proposition 7.4.2] works the same in the real case, and in particular:
@@
Remark. By splitting into the real and imaginary matrix parts, every complex real selfadjoint bounded multivalued nc function on $K_{c}$ is an $F+iG$. We may also define $c(F,G)$ in the obvious way. It is probably true by an adjustment of the proof of Proposition 4.2 that a complex bounded nc multivalued selfadjoint function $F+iG$ on $K_{c}$ is complex convex (resp. lowersemicontinuous) if and only if $c(F,G)$ is. We have not taken the time to check this.
@@
For $x\in K_{p}$ define $\tilde{F}(x)$ to be the intersection in the statement of the theorem. We have by [14, Theorem 7.4.3] that
@@
It seems that [14, Corollary 7.4.4] works the same in the real case. Finally (for Section 7.4), the noncommutative analogue of a result of Mokobodzki works with the same proof:
@@
Turning to [14, Section 7.5], this contains two remarkable results. The first says that the ‘convex envelope encodes information about the set of representing maps of a point’ [14]. Again the complex proof works.
@@
$\mu(\bar{F})=\cap_{m\in\mathbb{N}}\cap_{g\leq I_{m}\otimes F}\;\{\alpha\in(M_{n}(M_{p}(\mathbb{C}))_{\rm sa}:\mu(g)\leq I_{m}\otimes\alpha\},$
@@
Define $\tilde{\mu}$ to be the intersection in the statement of the theorem. We have by [14, Theorem 7.5.3] that

### v3

@@ 3. Complexification of extremals
and $\sum\,c(\alpha_{k})^{\mathsf{T}}c(\alpha_{k})=1_{2n}.$ Since $c(x)$ is pure in $K$, we have that a positive scalar multiple of $c(\alpha_{k})$ is a real isometry matrix $\eta_{k}\in M_{2n_{k},2n}(\mathbb{R})$, with $\eta_{k}^{\mathsf{T}}c(x_{k})\eta_{k}=c(x)$. It is then easy to see (since $c$ is one-to-one and ‘modular’) that a positive scalar multiple of $\alpha_{k}$ is a complex isometry $\beta_{k}\in M_{n_{k},n}(\mathbb{C})$, with $\beta_{k}^{*}x_{k}\beta_{k}=x$. So $x$ is pure in $K_{c}$. The rest follows since ‘extreme’ is ‘pure and maximal’. ∎
@@
The converse of the lemma is false even in the simplest example of $K$ singleton: $1$ is pure and extreme in $\mathrm{ncS}(\mathbb{C})$ but $c(1)=1\oplus 1$ is not pure, and is clearly not nc extreme. The class of extreme points considered in this lemma is interesting in some examples, but not in others.
@@
Remarks. (On real parts of complex matrix states.) 1) By the argument in [32, Proposition 5.3.3], every pure state $\varphi$ on a real operator system $V$ is the restriction to $V$ of the real part of a pure state on $V_{c}$ (not usually $\varphi_{c}$). However not every pure (resp. nc extreme) element $x\in K$ is the real part of a pure (resp. nc extreme) element of $K_{c}$, sadly. See [36], or 2) below, or Example 3.4. In the latter example, since $\ell^{\infty}_{2}$ has no irreducible representation into $M_{n}(\mathbb{C})$ for $n\geq 2$ we have no extreme points in level 2 (or higher) of the complexification. Similarly for pure points: in that example there are no pure elements of $K_{2}={\rm UCP}(\mathbb{C},M_{2}(\mathbb{C}))$ by 6.1.7 in [16].
@@
The classical argument that pure/extreme states are the real part of a pure state extension to the complexification is a standard one in convexity (finding an extreme point of the set of extensions and showing that this is extreme in the superspace). That the nc extension of this fails illustrates the failure in the nc situation of this very common trick in convexity. We do not know a characterization of when a nc pure (resp. nc extreme) in K is the real part of some pure (resp. nc extreme) in $K_{c}.$
@@
2) (A quaternionic counterexample.) Similarly the identity map on $M_{2}(\mathbb{C})$ is nc extreme, so maximal. However its real part $\psi$, which is Re, restricted to the copy of the quaternions $\mathbb{H}$, is not real maximal, so not extreme. This can be seen from Proposition 2.12 (2), or by the following direct argument: $\psi$ is the compression to the first two rows and columns of the usual (irreducible) representation of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. So it is not maximal by Proposition 2.12 (2). It is interesting that $\psi$ is pure, by Proposition 2.12 (1). (Cf. [36, Example 3.27].)
@@
The latter also answers the question: For a real ucp map $\varphi$, there exists a maximal dilation of its complexification. Is the real part of this maximal dilation a maximal dilation of $\varphi$? In the example above, the complexification of $\psi$ has maximal dilation $x\mapsto x\oplus x$. The real part of this is a real dilation of $\psi$. It is the map $\psi\oplus\psi$. However this is not maximal, since a summand of a maximal is easily seen to be maximal, and $\psi$ is not maximal. This also follows by Proposition 2.12 (2).
@@
For $K$ a complex compact nc convex set, if $x\in K$ is complex pure (resp. nc extreme, maximal) in $K$ then $x$ is real pure (resp. nc extreme, maximal) in $K$. Conversely, if $x$ is real nc extreme and complex-irreducible then $x$ is complex nc extreme and irreducible in $K$. In particular, in general
@@
For the next part we first note that we already showed in Lemma 3.1 and Corollary 2.7 that if $x\in K$ is complex maximal (resp. extreme) in $K_{c}$ then $x$ is complex maximal (resp. extreme) in $K$. Suppose that $x$ is real nc extreme in $K$ and complex-irreducible. Then $x$ is complex nc extreme in $K_{c}$ by Corollary 3.2. Then $x$ is real maximal in $K$, and hence complex maximal by Theorem 2.25. A variant of the argument in Corollary 3.2 shows that if $x$ is irreducible for $M_{n}(\mathbb{C})$ then it is complex extreme. Finally we give an example where $\partial_{\mathbb{C}}K\neq\partial_{\mathbb{R}}K$ in the next remark. ∎
Remark. By a slight variant of the argument for the last assertion of Corollary 3.2, $x\in K_{1}$ is real extreme in $K_{1}$ if and only if $x$ is complex extreme in $K$. However generally the converses in Lemma 3.8 fail for pure (resp. extreme). E.g. if $x$ is real nc extreme in $K$ then $x$ need not be complex extreme in $K$. This fails even in $K_{2}$. To see this consider an extreme $x\in K$ such that $x$ is not complex nc extreme in $K_{c}$, such as in Examples 3.4 or 3.6 (or Lemma 3.8). However by the Remark above Example 3.4, $x$ is real nc extreme in $K_{c}$. Similarly $x$ is real pure in $K_{c}$, but not complex pure there (if it were it would be complex extreme since it is also complex maximal).
@@ 4. Noncommutative convex functions
In the complex case this is Remark $7.2.2$ in [16] or [14, Proposition 16.13.2]. We include a proof, partly to serve as a model for the later more complicated multivalued nc function case. For $f\in M_{n}(B(K))$, we have that $f$ is graded, and so $\text{Epi}(f)$ is graded. Also $f$ preserves direct sums, which results in $\text{Epi}(f)$ being closed under direct sums. Indeed, let $(x_{i},\beta_{i})\in\text{Epi}_{m_{i}}(f)$, and let $\alpha_{i}\in M_{k,m_{i}}(\mathbb{R})$ be a family of isometries such that $\sum_{i}\alpha_{i}\alpha_{i}^{\mathsf{T}}=I_{k}$. We need to check that
@@
Remark. Similarly, the Hansen-Pedersen type application in [16, Proposition 7.2.3] holds in the real case, showing that $f$ is real nc convex if and only if it is classically convex at every ‘level’.
@@
Let $f\in M_{n}(B(K))_{\rm sa}$. Then $f$ is real nc convex (resp. lowersemicontinuous) if and only if $f_{c}=\Psi(f+i0)$ is complex convex (resp. lowersemicontinuous) as an element of $M_{n}(B(K_{c}))$.
@@
Suppose that $f+ig\in M_{n}(B(K))_{c}=M_{n}(B(K_{c}))$ for a real compact nc convex set $K$. Then $f+ig$ is complex nc convex if and only if $c(f,g)$ is real nc convex in $M_{2n}(B(K))$. If these hold then $f$ is real nc convex.
@@
Now $(f+ig)(x+iy)=u^{*}(f(c(x,y))+ig(c(x,y)))u$. Since $u(I_{n}\otimes\alpha)=(I_{2n}\otimes\alpha)u$, it suffices to show that
@@
Now [16, Proposition 7.2.8] is provable by complexification:
@@ 5. Convex multivalued functions and convex envelopes
We now extend our real versions of convexity and lower semicontinuity to bounded multivalued selfadjoint nc functions $F$. We define the latter in the real case as in [16, Definition 7.3.1] except of course matrices have real entries (including unitary matrices, naturally). Thus e.g. $F$ is bounded if there is a constant $\lambda>0$ such that for every $\beta\in F(x)$ there is an $\alpha\in F(x)$ with $\alpha\leq\beta$ and $\|\alpha\|\leq\lambda$. The graph ${\rm Graph}(F)$ is defined by
@@
Convexity and lower semicontinuity of $F$ are defined just as in [16, Definition 7.3.2], except all matrices have real entries. That is, $F$ is convex (resp. lowersemicontinuous) if ${\rm Graph}(F)$ is nc convex (resp. closed). We say that $F\leq G$ if $G(x)\subseteq F(x)$ for all $x\in K$. Every real (bounded) nc function $f:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ gives a (bounded) multivalued selfadjoint nc function $F(x)=[f(x),\infty)$, with Epi $(f)={\rm Graph}(F)$. So $f$ is convex if and only if $F$ is convex.
@@
For a real nc convex set $K$, every real bounded nc multivalued selfadjoint function $F:K\to M_{m}(\mathcal{M}(\mathbb{R}))$ has a unique complex bounded nc multivalued extension $F_{c}:K_{c}\to M_{m}(\mathcal{M}(\mathbb{C}))$, and this is selfadjoint. Similarly every real bounded nc selfadjoint multivalued function $F:K\to M_{m}(\mathcal{M}(\mathbb{C}))$ has a unique complex bounded selfadjoint nc multivalued extension $F_{c}:K_{c}\to M_{m}(\mathcal{M}(\mathbb{C}))$. Finally, $F\mapsto F_{c}$ is an order embedding: $F\leq G$ if and only if $F_{c}\leq G_{c}$.
@@
$F_{c}(\beta(x+iy)\beta^{*})$ $=u^{*}\,F(c(\beta)c(x+iy)c(\beta)^{\mathsf{T}})\,u$
$=u^{*}\,c(\beta)F(c(x+iy))c(\beta)^{\mathsf{T}}\,u$
$=\beta u^{*}\,F(c(x+iy))\,u\beta^{*}$
@@
which shows the unitary invariance of the extension. A similar proof shows the direct sum property. To see that $F_{c}(x+iy)=F_{c}(x+iy)+M_{m}(\mathcal{M}(\mathbb{C}))^{+}$ choose $\alpha\in M_{m}(\mathcal{M}(\mathbb{C}))^{+}$. Then $\alpha=u^{*}c(\alpha)u.$ For $z\in F(c(x+iy))$ we know that $z+c(\alpha)\in F(c(x+iy))$. Hence
@@
Thus $F_{c}(x+iy)=F_{c}(x+iy)+M_{m}(\mathcal{M}(\mathbb{C}))^{+}$.
@@
$(x+iy)\oplus(x-iy)=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}c(x+iy)\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}$.
@@
[…]
which equals
[…]
Here we first used that $G(x\oplus y)=G(x)\oplus G(y)$. The last equation is a $2n\times 2n$ matrix with top-right and bottom-left corners being 0. Comparing the top-left corners of the matrices in the above equation gives $G(x+iy)=u^{*}\,F(c(x+iy))\,u$ as desired.
@@
so taking the real part of both sides, and using that Re is unitarily equivariant (see Lemma 2.5 in [8]), we get that $\pi(G)$ is unitarily equivariant. For direct sums, take $x,y\in K$ and then $G(x)\oplus G(y)=G(x\oplus y)$. Also Re respects direct sums, so taking Re of both sides gives that $\pi(G)$ respects direct sums.
Fix $x\in K_{m}$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$. If $z+\alpha\in\pi(G)(x)+M_{n}(M_{m}(\mathbb{R}))^{+}$ for $z={\rm Re}(y)$ where $y\in G(x)$ and $\alpha\in M_{n}(M_{m}(\mathbb{R}))^{+}$, then $y+\alpha\in G(x)$ and so $\text{Re}(y)+\alpha\in\pi(G)(x)$. Thus $\pi(G)$ is upward directed. To show boundedness, let $x\in K$ and $\beta\in\pi(G)(x)$ so that $\beta=\text{Re}\,y$ for some $y\in G(x+i0)$. Because $G$ is bounded there exists a $\lambda>0$ and $\alpha\in G(x+i0)$ with $\alpha\leq y$ and $\|\alpha\|\leq\lambda$. Taking the Re of the inequality and using that Re is a contraction gives $\text{Re}\,\alpha\leq\text{Re}\,y=\beta$ and $\|\text{Re}\,\alpha\|\leq\lambda$. Thus, $\lambda$ is a bound for the function $\pi(G)$.
@@
The converse follows from the last lemma. Alternatively, suppose $F_{c}$ is convex and let $x\in K$ and $\alpha$ be an isometry. Then (2) holds for $F$ because $F(\alpha^{\mathsf{T}}x\alpha)=F_{c}(\alpha^{\mathsf{T}}x\alpha)$ and since (2) holds for $F_{c}$.
@@
The rest of [16, Section 7.3] is the same in the real case. Turning to [16, Section 7.4] we define the real convex envelope $\bar{F}$ of a real bounded selfadjoint multivalued nc function $F:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ to be the multivalued nc function $\bar{F}:K\to M_{n}(\mathcal{M}(\mathbb{R}))$ with Graph $(\bar{F})=\overline{{\rm ncconv}}({\rm Graph}(F)).$ It is easy to see that [16, Proposition 7.4.2] works the same in the real case, and in particular:
@@
Remark. By splitting into the real and imaginary matrix parts, every complex real selfadjoint bounded multivalued nc function on $K_{c}$ is an $F+iG$. We may also define $c(F,G)$ in the obvious way. It is probably true by an adjustment of the proof of Lemma 4.3 that a complex bounded nc multivalued selfadjoint function $F+iG$ on $K_{c}$ is complex convex (resp. lowersemicontinuous) if and only if $c(F,G)$ is. We have not taken the time to check this.
@@
For $x\in K_{p}$ define $\tilde{F}(x)$ to be the intersection in the statement of the theorem. We have by [16, Theorem 7.4.3] that
@@
It seems that [16, Corollary 7.4.4] works the same in the real case. Finally (for Section 7.4), the noncommutative analogue of a result of Mokobodzki works with the same proof:
@@
Turning to [16, Section 7.5], this contains two remarkable results. The first says that the ‘convex envelope encodes information about the set of representing maps of a point’ [16]. Again the complex proof works.
@@
$\mu(\bar{F})=\cap_{m\in\mathbb{N}}\cap_{g\leq I_{m}\otimes F}\;\{\alpha\in(M_{n}(M_{p}(\mathbb{R}))_{\rm sa}:\mu(g)\leq I_{m}\otimes\alpha\},$
@@
Define $\tilde{\mu}$ to be the intersection in the statement of the theorem. We have by [16, Theorem 7.5.3] that